# Magnetic-field-tuned randomness in inhomogeneous altermagnets

### v1

@@ I Introduction
A recent classification of collinear magnetism revealed the existence of three different types of magnetic states: ferromagnetism, antiferromagnetism, and altermagnetism [1, 2, 3]. These different magnetic ground states, which are formally defined in the absence of spin-orbit coupling using the concept of spin groups [3, 4, 5, 6, 7], are distinguished according to the type of space group operation which, when combined with time-reversal, leaves the ground state invariant. In ferromagnets (FM), there is no such operation and a uniform splitting between spin-up and spin-down bands emerges. Time-reversed antiferromagnetic (AFM) ground states, on the other hand, are related by a translation or inversion, resulting in a symmetry-protected Kramers spin degeneracy throughout the entire Brillouin zone. Finally, time-reversed altermagnetic (AM) ground states are related by any crystalline operation that is not translation or inversion, such as rotation, mirror reflection, or non-symmorphic operations like glides or screw rotations. The implication of this symmetry is a nodal $d$-wave, $g$-wave, or $i$-wave spin-split band structure, with spin degeneracy preserved only along certain high-symmetry momentum space planes (for a recent review, see [8]). These unusual properties have motivated extensive theoretical works on the interplay between altermagnetism and other electronic phenomena such as topology [9, 10, 11, 12, 13, 14, 15, 16], electronic correlations [17, 18, 19, 20, 21, 22], superconductivity [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], non-trivial responses [38, 39, 40, 41, 42, 43, 44, 45, 46], and multiferroics [47, 48, 49].
@@ II Random-field Ising model with a uniform transverse field
To proceed, we write down an effective low-energy model for the coupled AM-strain degrees of freedom. Since the AM order parameter is Ising like, we model it in terms of a transverse-field Ising model Hamiltonian:
@@
Here, $\tau_{i}^{z}$ and $\tau_{i}^{x}$ are Pauli matrices, where the thermal average $\langle\tau_{i}^{z}\rangle$ is the AM order parameter. $J$ is an effective AM interaction that sets the scale of the thermal AM transition (and should not be confused with an exchange interaction) whereas the transverse field $\Gamma$ is a non-thermal parameter (such as doping or pressure) that promotes quantum fluctuations, i.e., tunneling between local pseudo-spin states. Thus, as $\Gamma$ is enhanced, the ordered state is suppressed and the system is driven to a quantum critical point (QCP).
@@ VI Discussion and Conclusions
We emphasize that the tuning properties of the external magnetic field become much more pronounced if the system is closer to the AM-PM phase boundary, in which case the critical fields $H_{z,c}^{\pm}$ become small. Nevertheless, it is instructive to estimate the typical field magnitude $H_{\lambda}$ that sets the scale for many of the phenomena discussed here to occur. Minimizing the free energy $F(\Phi,\varepsilon)$ of piezomagnetically coupled order parameter and strain with respect to uniform strain $\varepsilon$ yields the linear response relation $\varepsilon=\lambda\Phi H_{z}/C_{0}$, which allows us to directly identify the piezomagnetic tensor element $\Lambda=\lambda\Phi/C_{0}$. Note that $\Lambda$ has dimensions of inverse magnetic field, since $\Phi$ is dimensionless, $C_{0}$ has dimensions of energy, and $\lambda$ has dimensions of energy times inverse field. Assuming $\Phi=\mathcal{O}(1)$ deep inside the AM phase, this enables us to relate the magnetic field scale $H_{\lambda}$ with other quantities:
$H_{\lambda}=\frac{\sqrt{C_{0}J}}{\lambda}\sim\frac{1}{\Lambda}\sqrt{\frac{J}{C_{0}}}$ (85)
Now, the energy scale associated with the shear modulus $C_{0}$ is typically of the order of $30$ eV, corresponding to an elastic constant of the order of $\mathcal{C}_{0}\sim 50$ GPa multiplied by a unit cell volume of the order of $v\sim 100$3, i.e., $C_{0}=\mathcal{C}_{0}v$. Conversely, the energy scale $J$ can be estimated from the typical AM transition temperature, giving $10$ meV, and thus $H_{\lambda}\sim 0.02/\Lambda$ As for the piezomagnetic coupling $\Lambda$, we consider rutiles like CoF2, FeF2, and MnF2, which are $d_{xy}$-wave altermagnets. The quantity reported in the literature for CoF2 [129, 130], $\bar{\Lambda}=6\times 10^{-3}\mu_{B}/\mathrm{GPa}$ , is actually the magnetization induced by applied stress, $m_{z}=\bar{\Lambda}\sigma_{xy}$. Since $m_{z}=\chi_{0}H_{z}$ and $\sigma_{xy}=\mathcal{C}_{0}\varepsilon$, we thus have $1/\Lambda=\bar{\Lambda}\mathcal{C}_{0}/\chi_{0}\sim 6T$. In the last step, we approximated the uniform magnetic susceptibility to a characteristic value for these rutiles [131, 132], $\chi_{0}\sim 0.05\mu_{B}/T$. Putting everything together, we find $H_{\lambda}\sim 0.1T$, which is an accessible value. We can also use these estimates to express the disorder strength in terms of the root mean square strain $\bar{\varepsilon}_{s}$, obtaining $W=\bar{\varepsilon}_{s}\sqrt{C_{0}/4J}\sim 30\bar{\varepsilon}_{s}$. This implies that $W\sim 1$, which ensures the existence of reentrant AM behavior, requires $\bar{\varepsilon}_{s}\sim 3\%$, which is a reasonable value.

### v2

@@ I Introduction
A recent classification of collinear magnetism revealed the existence of three different types of magnetic states: ferromagnetism, antiferromagnetism, and altermagnetism [1, 2, 3]. These different magnetic ground states, which are formally defined in the absence of spin-orbit coupling using the concept of spin groups [3, 4, 5, 6, 7], are distinguished according to the type of space group operation which, when combined with time-reversal, leaves the ground state invariant. In ferromagnets (FM), there is no such operation and a uniform splitting between spin-up and spin-down bands emerges. Time-reversed antiferromagnetic (AFM) ground states, on the other hand, are related by a translation or inversion, resulting in a symmetry-protected Kramers spin degeneracy throughout the entire Brillouin zone. Finally, time-reversed altermagnetic (AM) ground states are related by any crystalline operation that is not translation or inversion, such as rotation, mirror reflection, or non-symmorphic operations like glides or screw rotations. The implication of this symmetry is a nodal $d$-wave, $g$-wave, or $i$-wave spin-split band structure, with spin degeneracy preserved only along certain high-symmetry momentum space planes (for a recent review, see [8]). These unusual properties have motivated extensive theoretical works on the interplay between altermagnetism and other electronic phenomena such as topology [9, 10, 11, 12, 13, 14, 15, 16], electronic correlations [17, 18, 19, 20, 21, 22, 23], superconductivity [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38], non-trivial responses [39, 40, 41, 42, 43, 44, 45, 46, 47, 48], and multiferroics [49, 50, 51].
@@ II Random-field Ising model with a uniform transverse field
To proceed, we write down an effective low-energy model for the coupled AM-strain degrees of freedom. As explained above, for a tetragonal lattice, the $d$-wave AM order parameter transforms as either the $B_{1g}^{-}$ or the $B_{2g}^{-}$ irrep, which corresponds to an Ising-like scalar order parameter ($\Phi$ above). To account for the role of spatial and quantum fluctuations, we promote this homogeneous order parameter to a coarse-grained local Ising (i.e., pseudo-spin) variable $\tau^{z}_{i}$, such that $\langle\tau_{i}^{z}\rangle=\Phi$. Spatial fluctuations are encoded in the effective interaction between the pseudo-spins, whereas quantum fluctuations appear as a transverse field to the pseudo-spin. Thus, our phenomenological Hamiltonian for the AM phase is that of a transverse-field Ising model:
@@
where, as discussed above, $\tau_{i}^{z}$ and $\tau_{i}^{x}$ are Pauli matrices. $J$ is an effective AM interaction, whose origin depends on the microscopic mechanism behind AM order, that sets the scale of the thermal AM transition (and should not be confused with an exchange interaction). The transverse field $\Gamma$ is a non-thermal parameter (such as doping or pressure) that promotes quantum fluctuations, i.e., tunneling between local pseudo-spin states. Thus, as $\Gamma$ is enhanced, the ordered state is suppressed and the system is driven to a quantum critical point (QCP). Note that, being phenomenological, our approach is agnostic about the microscopic origin of the AM order parameter. For instance, in the Lieb lattice model of Ref. [12], $\tau_{i}^{z}$ could correspond to the staggered magnetization between the two sublattices related by a $90^{\circ}$ rotation, coarse-grained over a single square plaquette. It could also correspond to the average magnetic octupole moment on the third site of the Lieb lattice, or to some combination of them [43] determined by the energetics of the microscopic model [87]. The key point is that, regardless of the microscopic mechanism, the AM ordered state will occur due to the condensation of an Ising-like order parameter that transforms as the appropriate irrep.
@@ VI Discussion and Conclusions
We emphasize that the tuning properties of the external magnetic field become much more pronounced if the system is closer to the AM-PM phase boundary, in which case the critical fields $H_{z,c}^{\pm}$ become small. Nevertheless, it is instructive to estimate the typical field magnitude $H_{\lambda}$ that sets the scale for many of the phenomena discussed here to occur. Minimizing the free energy $F(\Phi,\varepsilon)$ of piezomagnetically coupled order parameter and strain with respect to uniform strain $\varepsilon$ yields the linear response relation $\varepsilon=\lambda\Phi H_{z}/C_{0}$, which allows us to directly identify the piezomagnetic tensor element $\Lambda^{-1}=\lambda\Phi/C_{0}$. Note that $\Lambda$ has dimensions of magnetic field, since $\Phi$ is dimensionless, $C_{0}$ has dimensions of energy, and $\lambda$ has dimensions of energy times inverse field. Assuming $\Phi=\mathcal{O}(1)$ deep inside the AM phase, this enables us to relate the magnetic field scale $H_{\lambda}$ with other quantities:
$H_{\lambda}=\frac{\sqrt{C_{0}J}}{\lambda}\sim\Lambda\sqrt{\frac{J}{C_{0}}}$ (85)
Now, the energy scale associated with the shear modulus $C_{0}$ is typically of the order of $30$ eV, corresponding to an elastic constant of the order of $\mathcal{C}_{0}\sim 50$ GPa multiplied by a unit cell volume of the order of $v\sim 100$3, i.e., $C_{0}=\mathcal{C}_{0}v$. Conversely, the energy scale $J$ can be estimated from the typical AM transition temperature, giving $10$ meV, and thus $H_{\lambda}\sim 0.02\Lambda$ As for the piezomagnetic coupling $\Lambda$, we consider rutiles like CoF2, FeF2, and MnF2, which are $d_{xy}$-wave altermagnets. The quantity reported in the literature for CoF2 [131, 132], $\bar{\Lambda}=6\times 10^{-3}\mu_{B}/\mathrm{GPa}$ , is actually the magnetization induced by applied stress, $m_{z}=\bar{\Lambda}\sigma_{xy}$. Since $m_{z}=\chi_{0}H_{z}$ and $\sigma_{xy}=\mathcal{C}_{0}\varepsilon$, we thus have $\Lambda=\bar{\Lambda}\mathcal{C}_{0}/\chi_{0}\sim 6T$. In the last step, we approximated the uniform magnetic susceptibility to a characteristic value for these rutiles [133, 134], $\chi_{0}\sim 0.05\mu_{B}/T$. Putting everything together, we find $H_{\lambda}\sim 0.1T$, which is an accessible value. We can also use these estimates to express the disorder strength in terms of the root mean square strain $\bar{\varepsilon}_{s}$, obtaining $W=\bar{\varepsilon}_{s}\sqrt{C_{0}/4J}\sim 30\bar{\varepsilon}_{s}$. This implies that $W\sim 1$, which ensures the existence of reentrant AM behavior, requires $\bar{\varepsilon}_{s}\sim 3\%$, which is a reasonable value. This estimate assumes that the most significant contribution to the FM fluctuations comes from the same degrees of freedom responsible for AM order. In contrast, the presence of separate FM degrees of freedom would suppress our estimate for $\lambda$. Whether this approximation applies to the rutiles requires further investigation.